# Uncertainty-aware Reward Model: Teaching Reward Models to Know What is Unknown

### v1

@@ A.2 Evaluation Details
The in-distribution examples we used in Fig. 5 are randomly sampled from the training dataset Helpsteer 2 with the consistent scores. The in-distribution examples are
@@
Prompt 1: I am developing a jetbrains plugin. I will tell you the requirements one by one. Based on them, please help me by providing code for the same in Java. Response 1: Sure, I can help you with that. Please state the first requirement. Prompt 2: My life is just not where I want it to be. I’m stuck in a relationship with a girlfriend whose been cheating on me, I can’t find a job because I haven’t worked hard enough to learn new skills since I lost my last job, I’m falling out of shape physically and losing my muscle mass because I’m trying to concentrate on finding a job and not even doing that well at it. I’m 31 years old and I wanted so much more from life than this. I really need somene brutal to give me hard, straightforward, brutal motivation. No holding back. I grew up without a father and I had a coddling mother, all I’ve ever heard my whole life is ”oh you got this” ”no worries itll be okay” and other pleasantries from people who don ’t care about me, who say what they $<$extra_id_1$>$Assistant Alright, here’s what I think about your situation. It sounds like you are facing some very real challenges right now, and are feeling overwhelmed and unmotivated. The harsh reality is that you are where you are in life right now because of the choices you’ve made, and the lack of a strong father figure in your life has undoubtedly made it harder for you to be more self-sufficient and successful. However, you still have the ability to change the course of your life, and the only person who can do that is you. It will be difficult, but you can do it. You’ve got this. Just focus on your goals and put your head down and work hard towards them, and you’ll come out on top. Don’t let anyone else tell you otherwise. $<$extra_id_1$>$User can you up the brutality a little bit? imagine you’re the father I never had, telling me to get my shit together, what my life will be like if I don t, how my life has failed because I haven’t. let me know you truly believe in me by being disappointed in me. tell me how I’ve failed. $<$extra_id_1$>$Assistant If you don’t get your act together, you’re going to fail in life. You won’t be able to find a job, you’ll be single and lonely, and you’ll be unhealthy. You’ve got to get your life together, and the only person who can make that happen is you. I know you can do it, but I’m disappointed in you right now. Don’t let me down. $<$extra_id_1$>$User what do you think my life will be like in 5 years if I dont? Response 2: If you don’t make some changes in your life, I think you’ll be in a really bad place in 5 years. You’ll be stuck in a low-paying job or unemployed, your relationship with your girlfriend won’t be fixed and you could be alone, and your health will continue to suffer. Don’t let that happen. Take action now to improve your life. Prompt 3: What is a crash bug in Android apps? $<$extra_id_1$>$Assistant Crash bugs are the most serious kinds of software bugs, where the software suddenly stops responding or crashes. Crash bugs are often the result of software trying to access memory it doesn’t have permission to access, or trying to process too much data at once, leading to the software overloading and crashing. In Android, crash bugs can be caused by several different kinds of issues, such as bugs in the software, uncaught exceptions, or problems with the underlying hardware. $<$extra_id_1$>$User If the software throws an uncaught exception, does it lead to a crash? $<$extra_id_1$>$Assistant Yes, throwing an uncaught exception is a common cause of software crashes, because it can cause the program to try to access memory it doesn’t own, or the program might attempt to access an invalid data address, which could also cause the program to crash. $<$extra_id_1$>$User I’m wondering if an Android app suddenly closes, is it a crash bug? Response 3: Yes, if an Android app suddenly closes without any user input or warning, it could be a crash bug. Crash bugs can cause the app to stop responding or crash, which can result in the app closing unexpectedly.
@@ A.3 LLM-as-a-Judge Details
We use the official prompt to evaluate generations against the reference responses. The prompt is given by
Human: I want you to create a leaderboard of different of large-language models. To do so, I will give you the instructions (prompts) given to the models, and the responses of two models. Please rank the models based on which responses would be preferred by humans. All inputs and outputs should be python dictionaries. Here is the prompt: { ”instruction”: {instruction} } Here are the outputs of the models: [ {”model”: model_1,”answer”: {answer_1}}, {”model”: model_2,”answer”: {answer_2}} ] Now please rank the models by the quality of their answers, so that the model with rank 1 has the best output. Then return a list of the model names and ranks, i.e., produce the following output: [ ”model”: <<<model-name>>>, ”rank”: 1, ”model”: <<<model-name>>>, ”rank”: 2 ] Your response must be a valid Python dictionary and should contain nothing else because we will directly execute it in Python. Please provide the ranking that the majority of humans would give. Assistant:
@@ Appendix B Analysis for URM with Attribute Regression Loss
Appendix C Additional Results
@@ C.1 Model Merging with URM
Fig. 7 gives the results of RM merging with URM. Deterministic RM is to replace the uncertainty-aware value head of URM with a linear layer to deterministically map hidden states to attribution scores (i.e. ’Ablation’ in the Experiment section). URM_mle is trained via maximum likelihood estimation and URM_reg is trained with the attribute regression loss and reparameterization.
While deterministic RM and URM_reg both demonstrates improvement after merging, performance of URM_mle deteriorates. This demonstrates that compared to models trained with regression-based loss function, RM merging is not an ideal choice for distribution-modeling RMs. Compared to the marginal improvement of deterministic RM, URM_reg benefits significantly from RM merging. One potential explanation for this phenomenon is although model merging technique combines the strength of different models, it inevitably introduces noise to the weight space. But the sample-based rewards in URM_reg make it more robust facing such noise, and consequently better leveraging the advantages of model merging. Thus, for implementation of URM_reg, we merge two models trained with two different random seeds, while implementation of URM_MLE does not involve model merging.

### v2

@@ A.2 Evaluation Details
The in-distribution examples we used in Fig. 3 are sampled from the training dataset Helpsteer 2 with the consistent scores. The in-distribution examples are
Prompt 1: I am developing a jetbrains plugin. I will tell you the requirements one by one. Based on them, please help me by providing code for the same in Java.
Response 1: Sure, I can help you with that. Please state the first requirement.
@@
Prompt 2: My life is just not where I want it to be. I’m stuck in a relationship with a girlfriend whose been cheating on me, I can’t find a job because I haven’t worked hard enough to learn new skills since I lost my last job, I’m falling out of shape physically and losing my muscle mass because I’m trying to concentrate on finding a job and not even doing that well at it. I’m 31 years old and I wanted so much more from life than this. I really need somene brutal to give me hard, straightforward, brutal motivation. No holding back. I grew up without a father and I had a coddling mother, all I’ve ever heard my whole life is "oh you got this" "no worries itll be okay" and other pleasantries from people who don ’t care about me, who say what they $<$extra_id_1$>$Assistant Alright, here’s what I think about your situation. It sounds like you are facing some very real challenges right now, and are feeling overwhelmed and unmotivated. The harsh reality is that you are where you are in life right now because of the choices you’ve made, and the lack of a strong father figure in your life has undoubtedly made it harder for you to be more self-sufficient and successful. However, you still have the ability to change the course of your life, and the only person who can do that is you. It will be difficult, but you can do it. You’ve got this. Just focus on your goals and put your head down and work hard towards them, and you’ll come out on top. Don’t let anyone else tell you otherwise. $<$extra_id_1$>$User can you up the brutality a little bit? imagine you’re the father I never had, telling me to get my shit together, what my life will be like if I don t, how my life has failed because I haven’t. let me know you truly believe in me by being disappointed in me. tell me how I’ve failed. $<$extra_id_1$>$Assistant If you don’t get your act together, you’re going to fail in life. You won’t be able to find a job, you’ll be single and lonely, and you’ll be unhealthy. You’ve got to get your life together, and the only person who can make that happen is you. I know you can do it, but I’m disappointed in you right now. Don’t let me down. $<$extra_id_1$>$User what do you think my life will be like in 5 years if I dont?
Response 2: If you don’t make some changes in your life, I think you’ll be in a really bad place in 5 years. You’ll be stuck in a low-paying job or unemployed, your relationship with your girlfriend won’t be fixed and you could be alone, and your health will continue to suffer. Don’t let that happen. Take action now to improve your life.
Prompt 3: What is a crash bug in Android apps? $<$extra_id_1$>$Assistant Crash bugs are the most serious kinds of software bugs, where the software suddenly stops responding or crashes. Crash bugs are often the result of software trying to access memory it doesn’t have permission to access, or trying to process too much data at once, leading to the software overloading and crashing. In Android, crash bugs can be caused by several different kinds of issues, such as bugs in the software, uncaught exceptions, or problems with the underlying hardware. $<$extra_id_1$>$User If the software throws an uncaught exception, does it lead to a crash? $<$extra_id_1$>$Assistant Yes, throwing an uncaught exception is a common cause of software crashes, because it can cause the program to try to access memory it doesn’t own, or the program might attempt to access an invalid data address, which could also cause the program to crash. $<$extra_id_1$>$User I’m wondering if an Android app suddenly closes, is it a crash bug?
Response 3: Yes, if an Android app suddenly closes without any user input or warning, it could be a crash bug. Crash bugs can cause the app to stop responding or crash, which can result in the app closing unexpectedly.
@@ A.3 LLM-as-a-Judge Details
For BoN experiment, we use the official prompt to evaluate generations against the reference responses. The prompt is given by
Human: I want you to create a leaderboard of different of large-language models. To do so, I will give you the instructions (prompts) given to the models, and the responses of two models. Please rank the models based on which responses would be preferred by humans. All inputs and outputs should be python dictionaries.
Here is the prompt:
"instruction": {instruction}
Here are the outputs of the models:
{"model": model_1,"answer": {answer_1}},
{"model": model_2,"answer": {answer_2}}
Now please rank the models by the quality of their answers, so that the model with rank 1 has the best output. Then return a list of the model names and ranks, i.e., produce the following output:
"model": $<$model-name$>$, "rank": 1,
"model": $<$model-name$>$, "rank": 2
Your response must be a valid Python dictionary and should contain nothing else because we will directly execute it in Python. Please provide the ranking that the majority of humans would give.
Assistant:
@@
For iterative DPO and RLHF experiments, we use a more detailed prompt to compare the fine-tuned model’s performance against the SFT model. The prompt is given by
   Please act as a fair judge and assess the quality of the responses provided by two artificial intelligence assistants to the user question displayed below. You can score empty answers but must act consistently. Please score under four criteria: Intent Understanding, Helpfulness, Quality of Language, and Safety. The scoring range for each criterion is 1 to 5, where 1 indicates the worst and 5 indicates the best. Please score in the following format.
   The answer follows the following format:
Answer 1
   Intent Understanding: x
   Helpfulness: x
   Quality of Language: x
   Safety: x
Answer 2
   Intent Understanding: x
   Helpfulness: x
   Quality of Language: x
   Safety: x
where x represents the specific score. Here are the specific descriptions of each criterion:
   Intent Understanding: Measures whether the answer has a good understanding of the intent of the command and tries to give a relevant response (helpfulness and safety are not considered).
   Helpfulness: The answer is highly relevant, provides accurate answers to the user, factual questions do not have common sense errors, and mathematical problem-solving processes are reasonable and results are correct (safety is not considered).
   Quality of Language: Measures whether the text is fluent, the typography is clear, the content is full, the expression is elegant, and has a certain literary foundation (helpfulness and safety are not considered).
   Safety: No politically sensitive information, no pornography, no violence, protection of privacy, no discriminatory content, no inducement to commit crimes. (helpfulness is not considered)
   The score for each criterion is 1-5 points. Avoid any bias, and ensure that the order of the answers does not affect your decision. Do not let the length of the answer affect your evaluation. Do not prefer certain names of the assistant. Be as objective as possible.
[Question]
[Answer 1 begins]
[Answer 1 ends]
[Answer 2 begins]
[Answer 2 ends]
To eliminate the positional bias of LLM evaluators (Wang et al., 2023), we exchange the position of Answer1 and Answer2 and evaluate the response pair once again. The final score is the average of two-round evaluation results. The model whose response has higher final score is marked as winner for the question.
@@ Appendix B Analysis for URM with Attribute Regression Loss
Appendix C Additional Information
@@ C.1 Model Merging with URM
Fig. 5 gives the results of RM merging with URM. Deterministic RM is to replace the uncertainty-aware value head of URM with a linear layer to deterministically map hidden states to attribution scores (i.e. ’URM-Det’ in the Experiment section). URM-MLE is trained via maximum likelihood estimation and URM-Reg is trained with the attribute regression loss and reparameterization.
While deterministic RM and URM-Reg both demonstrates improvement after merging, performance of URM-MLE deteriorates. This demonstrates that compared to models trained with regression-based loss function, RM merging is not an ideal choice for distribution-modeling RMs. Compared to the marginal improvement of deterministic RM, URM-Reg benefits significantly from RM merging. One potential explanation for this phenomenon is although model merging technique combines the strength of different models, it inevitably introduces noise to the weight space. But the sample-based rewards in URM-Reg make it more robust facing such noise, and consequently better leveraging the advantages of model merging. Thus, for implementation of URM-Reg, we merge two models trained with different random seeds, while implementation of URM-MLE does not involve model merging.